(3) Activity Quiz. Please explain how to find that last two parts.


Our goal is to use this video to measure the accel
Answer #1

Solution:

The speed after 27 s after launch is calculated using the basic kinematic equation vf = vi +at

here vi=0 (that is , the shuttle started off from rest)

so after 27 s its velocity vf = 0+(6.14)(27 = 165.78 m/s

Its altitude at 27s :

y = vit+1/2gt^2

   = 1/2 * 9.8 * (27)2 = 3572.1 m = 3.57 km